OpenStudy (anonymous):

Show all work. The area of a rectangular swimming pool is given by 2x^2 + x - 10 ft^2 . One side length of the pool is given by 2x + 5 feet. What is an algebraic expression for the other side length of the pool? Simplify this, and include correct units as part of your answer.

OpenStudy (anonymous):

Area / length is the other side (2x^2 +x -10)/ (2x +5) 2x^2 +x -10 = 2x^2 +5x -4x -10 = x (2x +5) -2(2x +5) = (x-2)(2x +5) (2x^2 +x -10)/ (2x +5) = (x-2) (2x +5) / (2x +5) = x-2 Other side is x -2

OpenStudy (anonymous):

Thanks Rita!